OpenStudy (anonymous):

Determine algebraically whether f(x)=1+x+x^2 is even, odd or neither

OpenStudy (cwrw238):

for even function: f(x) = f(-x) for odd: -f(x) = f(-x)

OpenStudy (cwrw238):

f(x) = 1+ x + x^2 replace x by -x: f(-x) = 1 - x + (-x)^2 which shows its not even now work out -f(x) and see what you get